### 13 Feb 2024 v1

@@ Appendix D Correlation-based Transductive Learning
$=\frac{1}{2|\mathcal{A}|}\sum_{\boldsymbol{x^{\prime}}\in\mathcal{A}}\mathrm{I}\!\left(f_{\boldsymbol{x^{\prime}}};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right)$
${=}-\sum_{\boldsymbol{x^{\prime}}\in\mathcal{A}}\log\left(1-\mathrm{Cor}{}\left[f_{\boldsymbol{x^{\prime}}},y_{\boldsymbol{x}}\mid\mathcal{D}_{n-1}\right]^{2}\right)$
@@
where $(i)$ is detailed in example 8.5.1 of (Cover, 1999); and $(iii)$ follows from Jensen’s inequality and exponentiating. $(i)$ is a loose approximation which highlights that ITL takes into account the correlations between points in $\mathcal{A}$ while CTL dooes not.
@@ E.1 Stochastic Target Spaces
With this result, the previously derived convergence guarantees follow also for the stochastic target space setting.
@@ Appendix G Additional NN Experiments & Details
      compute embedding matrix $\widetilde{\boldsymbol{\Phi}}$ over domain $[S,A^{\prime}]$
@@
         compute acquisition function w.r.t. $A^{\prime}$, based on $\widetilde{\boldsymbol{\Phi}}$
@@
         update conditional embedding matrix $\widetilde{\boldsymbol{\Phi}}$
@@
In Section G.1, we detail metrics and hyperparameters. We describe in Sections G.2 and G.3 how to compute the (initial) conditional embedding matrix $\widetilde{\boldsymbol{\Phi}}$, and in Section G.4 how to update this matrix $\widetilde{\boldsymbol{\Phi}}$ to obtain conditional embeddings for batch selection.
@@ G.1 Experiment Details
Table 1: Hyperparameter summary of NN experiments.
@@
# of epochs (*) [^15] We train until convergence with respect to oracle validation accuracy. $5$
@@
We provide an overview of the hyperparameters used in our NN experiments in Table 1. The effect of noise standard deviation $\rho$ is small for all tested $\rho\in[1,100]$ (cf. ablation study in Table 2). [^16] We use a larger noise standard deviation $\rho$ in CIFAR-100 to stabilize the numerics of batch selection via conditional embeddings (cf. Table 2). $M$ denotes the size of the sample ${A\sim\mathcal{P}_{\!\!\mathcal{A}}}$. In each iteration, we select the target space $\mathcal{A}\leftarrow A^{\prime}$ as a random subset of $m$ points from $A$ [^17] This appears to improve the training, likely because it prevents overfitting to peculiarities in the finite sample $A$. We observe that choosing $m$ to match the number of classes yields good performance.
During each iteration, we select the batch $B$ by choosing the top-$b$ points according to the decision rule from a random sample from $\mathcal{P}_{\!\mathcal{S}}$ of size $k$ [^18] In large-scale problems, the work of Coleman et al. (2022) suggests to use an (approximate) nearest neighbor search to select the (large) candidate set rather than sampling u.a.r. from $\mathcal{P}_{\!\mathcal{S}}$. This can be a viable alternative to simply increasing $k$ and suggests future work.
@@
We use the ADAM optimizer (Kingma & Ba, 2014). In our CIFAR-100 experiments, we use a pre-trained EfficientNet-B0 (Tan & Le, 2019) model, and fine-tune the final and penultimate layers. We freeze earlier layers to prevent overfitting to the few-shot training data.
@@
instead of Equation 1, which amounts to adding $\rho^{2}$ to the diagonal of the covariance matrix before inversion. This appears to improve numerical stability, especially when using gradient embeddings. [^20] In our experiments, we observe that the effect various choices of $\rho$ on this slight adaptation of the ITL decision rule has negligible impact on performance. The more prominent effect of $\rho$ appears to arise from the batch selection via conditional embeddings (cf. Table 2).
G.2 Gradient vs. Latent Embeddings
@@
Figure 4: Comparison of gradient (“G-”) and latent embeddings (“L-”).
Using a neural network to parameterize $f$, we evaluate the following canonical approximations of $f$ by a stochastic process:
@@
    Latent Embeddings: Taking the latent representation $\boldsymbol{\phi}(\boldsymbol{x})$ from the penultimate layer and collecting them as rows in the design matrix $\boldsymbol{\Phi}$ of inputs $X$, we can view the last layer as the linear function ${\boldsymbol{f}_{\!\!X}=\boldsymbol{\Phi}\boldsymbol{\theta}}$. Approximating the weights by ${\boldsymbol{\theta}\sim\mathcal{N}({}\boldsymbol{\mu},\boldsymbol{\Sigma})}$ implies $\boldsymbol{f}_{\!\!X}\sim\mathcal{N}({}\boldsymbol{\Phi}\boldsymbol{\mu},\boldsymbol{\Phi}\boldsymbol{\Sigma}\boldsymbol{\Phi}^{\top})$. Interpreting the early layers as a feature encoder, thus uses the low-dimensional feature map akin to random feature methods (Rahimi & Recht, 2007).
@@
    Gradient Embeddings: Alternatively, one can approximate ${\boldsymbol{f}_{\!\!X}\sim\mathcal{N}({}\boldsymbol{\mu},\boldsymbol{J}\boldsymbol{\Sigma}\boldsymbol{J}^{\top})}$ where $\boldsymbol{J}=\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\boldsymbol{f}(X;\boldsymbol{\theta})$ is the Jacobian of NN parameters $\boldsymbol{\theta}$, effectively taking the gradient embeddings $\boldsymbol{g}(\boldsymbol{x})=\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,f(\boldsymbol{x};\boldsymbol{\theta})$ as feature representations (Khan et al., 2019; Holzmüller et al., 2023). This kernel is called the (empirical) neural tangent kernel (NTK) and the posterior mean of this GP approximates wide NNs trained with gradient descent (Arora et al., 2019). Kassraie & Krause (2022) derive bounds of $\gamma_{n}$ under this kernel.
    Rather than the NTK, we use in our experiments the gradient embeddings
    $\boldsymbol{g}(\boldsymbol{x})=\left.\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta^{\prime}}}\,\ell_{\mathrm{CE}}(f(\boldsymbol{x};\boldsymbol{\theta}),\hat{y}(\boldsymbol{x}))\right\rvert_{\boldsymbol{\theta}=\boldsymbol{\widehat{\theta}}}$
    where $\boldsymbol{\theta^{\prime}}$ refers to the parameters of the final output layer, $\boldsymbol{\widehat{\theta}}$ are the current parameter estimates, $\smash{\hat{y}(\boldsymbol{x})=\operatorname*{arg\,max}_{i}f_{i}(\boldsymbol{x};\boldsymbol{\widehat{\theta}})}$ are the associated predicted labels, and $\ell_{\mathrm{CE}}$ denotes the cross-entropy loss. This gradient embedding captures the potential update direction upon observing a new point (Ash et al., 2020). Moreover, Ash et al. (2020) show that for most neural networks, the norm of these gradient embeddings are a conservative lower bound to the norm assumed by taking any other proxy label $\hat{y}(\boldsymbol{x})$.
In the following, we denote the design matrix of inputs $X$ w.l.o.g. by $\boldsymbol{\Phi}$ (collecting the embeddings of all $\boldsymbol{x}\in X$ as rows), and the covariance matrix of $\boldsymbol{f}_{\!\!X}$ by $\widetilde{\boldsymbol{\Phi}}\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\boldsymbol{\Phi}\boldsymbol{\Sigma}\boldsymbol{\Phi}^{\top}$.
Unless otherwise specified, we use gradient embeddings in our experiments. We compare latent and gradient embeddings in Figure 4 and observe only negligible differences in performance.
G.3 Towards Quantifying Epistemic Uncertainty
@@
Figure 5: Comparison uncertainty quantification (i.e., estimation of $\boldsymbol{\Sigma}$) via a Laplace approximation (LA, Daxberger et al. (2021)) using a Kronecker factored log-likelihood Hessian approximation (Martens & Grosse, 2015). The results are shown for the MNIST experiment. We do not observe a performance improvement beyond the trivial approximation $\boldsymbol{\Sigma}=\boldsymbol{I}$.
@@
The poor performance of UnSa (cf. Section G.5) with this approximation suggests that with more sophisticated models, the performance of ITL and CTL can be further improved. Various approximations of $\boldsymbol{\Sigma}$ based on Laplace approximation (Daxberger et al., 2021), variational inference (Blundell et al., 2015), or Markov chain Monte Carlo (Maddox et al., 2019) have been studied. We also evaluate Laplace approximation (LA, Daxberger et al. (2021)) for estimating $\boldsymbol{\Sigma}$ but see no improvement (cf. Figure 5). Nevertheless, we believe that uncertainty quantification is a promising direction for future work, with the potential to improve performance of ITL and its variations substantially.
@@ G.4 Batch Selection via Conditional Embeddings
We will refer to Equation 5 as BaCE, short for batch selection via conditional embeddings. BaCE can be implemented efficiently using the Gaussian approximation of $\boldsymbol{f}_{\!\!X}$ from Section G.2 by iteratively conditioning on the previously selected points $\boldsymbol{x}_{n,1:i-1}$, and updating the embedding matrix $\smash{\widetilde{\boldsymbol{\Phi}}}$ using the closed-form formula for the variance of conditional Gaussians:
$\widetilde{\boldsymbol{\Phi}}\leftarrow\widetilde{\boldsymbol{\Phi}}-\frac{1}{\widetilde{\boldsymbol{\Phi}}_{j,j}+\rho^{2}}\widetilde{\boldsymbol{\Phi}}_{X,j}\widetilde{\boldsymbol{\Phi}}_{j,X}$ (25)
where $j$ denotes the index of the selected $\boldsymbol{x}_{n,i}$ within $X$ and $\rho^{2}$ is the noise variance. Note that $\smash{\widetilde{\boldsymbol{\Phi}}}_{j,j}$ is a scalar and $\smash{\widetilde{\boldsymbol{\Phi}}}_{X,j}$ is a row vector, and hence, this iterative update can be implemented efficiently.
We remark that Equations 4 and 5 are natural extensions of previous non-adaptive active learning methods, which typically maximize some notion of “distance” between points in the batch, to the “directed” setting (Ash et al., 2020; Zanette et al., 2021; Holzmüller et al., 2023; Pacchiano et al., 2024). BaCE simultaneously maximizes “distance” between points in a batch and minimizes “distance” to points in $\mathcal{A}$.
@@ G.5 Baselines
Figure 6: Comparison of baselines for the experiment of Figure 2.
@@
We compare to the following commonly used heuristic “undirected” decision rules:
  • BADGE (Ash et al., 2020) which selects a batch according to the kmeans++ initialization scheme, using the gradient embeddings $\boldsymbol{g}(\boldsymbol{x})$ as features.
  • MaxMargin (Settles & Craven, 2008):
    $\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}p_{1}(\boldsymbol{x};\boldsymbol{\widehat{\theta}}_{n-1})-p_{2}(\boldsymbol{x};\boldsymbol{\widehat{\theta}}_{n-1})$
    where $p_{1}$ and $p_{2}$ are the two largest class probabilities.
@@
  • Uncertainty sampling (UnSa) which with embeddings $\boldsymbol{\phi}_{n-1}$ corresponds to:
@@
    We evaluate UnSa with gradient embeddings. This acquisition function is similar to previously used “embedding length” or “gradient length” heuristics (Settles & Craven, 2008).
We summarize the results in Figure 6. We observe that most “undirected” decision rules perform worse (and often significantly so) than Random. This is likely due to frequently selecting points from the support of $\mathcal{P}_{\!\mathcal{S}}$ which are not in the support of $\mathcal{P}_{\!\!\mathcal{A}}$ since the points are “adversarial examples” that the model $\boldsymbol{\widehat{\theta}}$ is not trained to perform well on. In the case of MNIST, this can also partially be attributed to the well-known “cold-start problem” (Gao et al., 2020).
We also compare to the following “directed” decision rules:
@@
    $\angle_{\boldsymbol{\phi}}(\boldsymbol{x},\mathcal{A})\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\sum_{\boldsymbol{x^{\prime}}\in\mathcal{A}}\frac{\boldsymbol{\phi}(\boldsymbol{x})^{\top}\boldsymbol{\phi}(\boldsymbol{x^{\prime}})}{\left\|\boldsymbol{\phi}(\boldsymbol{x})\right\|_{2}\left\|\boldsymbol{\phi}(\boldsymbol{x^{\prime}})\right\|_{2}}.$
@@
    $\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\mathrm{H}[p(\boldsymbol{x};\boldsymbol{\widehat{\theta}}_{n-1})]\cdot\left(\frac{1}{|\mathcal{A}|}\angle_{\boldsymbol{\phi}_{n-1}}(\boldsymbol{x},\mathcal{A})\right)^{\beta}$
@@
We evaluate both baselines using latent embeddings (cf. Section G.2), as is typical.

### 12 Mar 2024 v2

@@ Appendix D Correlation-based Transductive Learning
${=}-\frac{1}{2|\mathcal{A}|}\sum_{\boldsymbol{x^{\prime}}\in\mathcal{A}}\log\left(1-\mathrm{Cor}{}\left[f_{\boldsymbol{x^{\prime}}},y_{\boldsymbol{x}}\mid\mathcal{D}_{n-1}\right]^{2}\right)$
@@
where $(ii)$ is detailed in example 8.5.1 of (Cover, 1999); and $(iii)$ follows from Jensen’s inequality and exponentiating. $(i)$ is a loose approximation which highlights that ITL takes into account the correlations between points in $\mathcal{A}$ while CTL dooes not.
@@ E.1 Stochastic Target Spaces
With this result, the previously derived convergence guarantees follow also for the stochastic target space setting. We leave a tighter analysis of stochastic target spaces to future work.
@@ Appendix G Additional NN Experiments & Details
      compute kernel matrix $\boldsymbol{K}$ over domain $[S,A^{\prime}]$
@@
         compute acquisition function w.r.t. $A^{\prime}$, based on $\boldsymbol{K}$
@@
         update conditional kernel matrix $\boldsymbol{K}$
@@
In Section G.1, we detail metrics and hyperparameters. We describe in Sections G.2 and G.3 how to compute the (initial) conditional kernel matrix $\boldsymbol{K}$, and in Section G.4 how to update this matrix $\boldsymbol{K}$ to obtain conditional embeddings for batch selection.
@@
Hübotter et al. (2024) provides additional comparisons with respect to a variance-based trasnductive decision rule.
@@ G.1 Experiment Details
Table 1: Hyperparameter summary of NN experiments. (*) we train until convergence on oracle validation accuracy.
@@
# of epochs (*) $5$
@@
We provide an overview of the hyperparameters used in our NN experiments in Table 1. The effect of noise standard deviation $\rho$ is small for all tested $\rho\in[1,100]$ (cf. ablation study in Table 2). [^17] We use a larger noise standard deviation $\rho$ in CIFAR-100 to stabilize the numerics of batch selection via conditional embeddings (cf. Table 2). $M$ denotes the size of the sample ${A\sim\mathcal{P}_{\!\!\mathcal{A}}}$. In each iteration, we select the target space $\mathcal{A}\leftarrow A^{\prime}$ as a random subset of $m$ points from $A$ [^18] This appears to improve the training, likely because it prevents overfitting to peculiarities in the finite sample $A$ (cf. Figure 10). We provide an ablation over $m$ in Section G.6.
During each iteration, we select the batch $B$ according to the decision rule from a random sample from $\mathcal{P}_{\!\mathcal{S}}$ of size $k$ [^19] In large-scale problems, the work of Coleman et al. (2022) suggests to use an (approximate) nearest neighbor search to select the (large) candidate set rather than sampling u.a.r. from $\mathcal{P}_{\!\mathcal{S}}$. This can be a viable alternative to simply increasing $k$ and suggests future work.
@@
We use the ADAM optimizer (Kingma & Ba, 2014). In our CIFAR-100 experiments, we use a pre-trained EfficientNet-B0 (Tan & Le, 2019), and fine-tune the final and penultimate layers. We freeze earlier layers to prevent overfitting to the few-shot training data.
@@
instead of Equation 1, which amounts to adding $\rho^{2}$ to the diagonal of the covariance matrix before inversion. This appears to improve numerical stability, especially when using gradient embeddings. [^21] In our experiments, we observe that the effect of various choices of $\rho$ on this slight adaptation of the ITL decision rule has negligible impact on performance. The more prominent effect of $\rho$ appears to arise from the batch selection via conditional embeddings (cf. Table 2).
G.2 Embeddings and Kernels
@@
Figure 4: Comparison of loss gradient (“G-”) and last-layer embeddings (“L-”).
Using a neural network to parameterize $f$, we evaluate the canonical approximations of $f$ by a stochastic process in the following.
An embedding $\boldsymbol{\phi}(\boldsymbol{x})$ is a latent representation of an input $\boldsymbol{x}$. Collecting the embeddings as rows in the design matrix $\boldsymbol{\Phi}$ of a set of inputs $X$, one can approximate the network by the linear function ${\boldsymbol{f}_{\!\!X}=\boldsymbol{\Phi}\boldsymbol{\beta}}$ with weights $\boldsymbol{\beta}$. Approximating the weights by ${\boldsymbol{\beta}\sim\mathcal{N}({}\boldsymbol{\mu},\boldsymbol{\Sigma})}$ implies that ${\boldsymbol{f}_{\!\!X}\sim\mathcal{N}({}\boldsymbol{\Phi}\boldsymbol{\mu},\boldsymbol{\Phi}\boldsymbol{\Sigma}\boldsymbol{\Phi}^{\top})}$. The covariance matrix ${\boldsymbol{K}_{\!XX}=\boldsymbol{\Phi}\boldsymbol{\Sigma}\boldsymbol{\Phi}^{\top}}$ can be succinctly represented in terms of its associated kernel ${k(\boldsymbol{x},\boldsymbol{x^{\prime}})=\boldsymbol{\phi}(\boldsymbol{x})^{\top}\boldsymbol{\Sigma}\boldsymbol{\phi}(\boldsymbol{x^{\prime}})}$. Here,
  • $\boldsymbol{\phi}(\boldsymbol{x})$ is the latent representation of $\boldsymbol{x}$, and
  • $\boldsymbol{\Sigma}$ captures the dependencies in the latent space.
While any choice of embedding $\boldsymbol{\phi}$ is possible, the following are common choices:
@@
    Last-Layer: A common choice for $\boldsymbol{\phi}(\boldsymbol{x})$ is the representation of $\boldsymbol{x}$ from the penultimate layer of the neural network (Holzmüller et al., 2023). Interpreting the early layers as a feature encoder, this uses the low-dimensional feature map akin to random feature methods (Rahimi & Recht, 2007).
@@
    Output Gradients (empirical NTK): Another common choice is $\boldsymbol{\phi}(\boldsymbol{x})=\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\boldsymbol{f}(\boldsymbol{x};\boldsymbol{\theta})$ where $\boldsymbol{\theta}$ are the network parameters (Holzmüller et al., 2023). Its associated kernel is known as the (empirical) neural tangent kernel (NTK) and the posterior mean of this GP approximates wide NNs trained with gradient descent (Arora et al., 2019; Khan et al., 2019; He et al., 2020). Kassraie & Krause (2022) derive bounds of $\gamma_{n}$ under this kernel. If $\boldsymbol{\theta}$ is restricted to the weights of the final linear layer, then this embedding is simply the last-layer embedding.
  3. 3.
    Loss Gradients: Another possible choice is
    $\boldsymbol{\phi}(\boldsymbol{x})=\left.\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\ell(\boldsymbol{f}(\boldsymbol{x};\boldsymbol{\theta}),\hat{y}(\boldsymbol{x}))\right\rvert_{\boldsymbol{\theta}=\boldsymbol{\widehat{\theta}}}$
    where $\ell$ is a loss function, $\hat{y}(\boldsymbol{x})$ is the predicted label, and $\boldsymbol{\widehat{\theta}}$ are the current parameter estimates Ash et al. (2020).
  4. 4.
    Outputs (empirical NNGP): Another possible choice is $\boldsymbol{\phi}(\boldsymbol{x})=\boldsymbol{f}(\boldsymbol{x})$, i.e., the output of the network. Its associated kernel is known as the (empirical) neural network Gaussian process (NNGP) kernel (Lee et al., 2018).
In the additional experiments from this appendix we use last-layer embeddings unless noted otherwise. We compare the performance of last-layer and the loss gradient embedding
$\boldsymbol{\phi}(\boldsymbol{x})=\left.\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta^{\prime}}}\,\ell_{\mathrm{CE}}(\boldsymbol{f}(\boldsymbol{x};\boldsymbol{\theta}),\hat{y}(\boldsymbol{x}))\right\rvert_{\boldsymbol{\theta}=\boldsymbol{\widehat{\theta}}}$ (26)
where $\boldsymbol{\theta^{\prime}}$ are the parameters of the final output layer, $\smash{\boldsymbol{\widehat{\theta}}}$ are the current parameter estimates, $\smash{\hat{y}(\boldsymbol{x})=\operatorname*{arg\,max}_{i}f_{i}(\boldsymbol{x};\boldsymbol{\widehat{\theta}})}$ are the associated predicted labels, and $\ell_{\mathrm{CE}}$ denotes the cross-entropy loss. This gradient embedding captures the potential update direction upon observing a new point (Ash et al., 2020). Moreover, Ash et al. (2020) show that for most neural networks, the norm of these gradient embeddings are a conservative lower bound to the norm assumed by taking any other proxy label $\hat{y}(\boldsymbol{x})$. In Figure 4, we observe only negligible differences in performance between this and the last-layer embedding.
G.3 Towards Uncertainty Quantification in Latent Space
@@
Figure 5: Uncertainty quantification (i.e., estimation of $\boldsymbol{\Sigma}$) via a Laplace approximation (LA, Daxberger et al. (2021)) over last-layer weights using a Kronecker factored log-likelihood Hessian approximation (Martens & Grosse, 2015) and the loss gradient embeddings from Equation 26. The results are shown for the MNIST experiment. We do not observe a performance improvement beyond the trivial approximation $\boldsymbol{\Sigma}=\boldsymbol{I}$.
@@
The poor performance of UnSa (cf. Section G.5) with this approximation suggests that with more sophisticated approximations, the performance of ITL and CTL can be further improved. Further research is needed to study the effect of more sophisticated approximations of “uncertainty” in the latent space. For example, with parameter gradient embeddings, the latent space is the network parameter space where various approximations of $\boldsymbol{\Sigma}$ based on Laplace approximation (Daxberger et al., 2021; Antorán et al., 2022), variational inference (Blundell et al., 2015), or Markov chain Monte Carlo (Maddox et al., 2019) have been studied. We also evaluate Laplace approximation (LA, Daxberger et al. (2021)) for estimating $\boldsymbol{\Sigma}$ but see no improvement (cf. Figure 5). Nevertheless, we believe that uncertainty quantification is a promising direction for future work, with the potential to improve performance of ITL and its variations substantially.
@@ G.4 Batch Selection via Conditional Embeddings
We will refer to Equation 6 as BaCE, short for batch selection via conditional embeddings. BaCE can be implemented efficiently using the Gaussian approximation of $\boldsymbol{f}_{\!\!X}$ from Section G.2 by iteratively conditioning on the previously selected points $\boldsymbol{x}_{n,1:i-1}$, and updating the kernel matrix $\boldsymbol{K}_{\!XX}$ using the closed-form formula for the variance of conditional Gaussians:
$\boldsymbol{K}_{\!XX}\leftarrow\boldsymbol{K}_{\!XX}-\frac{1}{\boldsymbol{K}_{\!\boldsymbol{x}_{j}\boldsymbol{x}_{j}}+\rho^{2}}\boldsymbol{K}_{\!X\boldsymbol{x}_{j}}\boldsymbol{K}_{\!\boldsymbol{x}_{j}X}$ (27)
where $j$ denotes the index of the selected $\boldsymbol{x}_{n,i}$ within $X$ and $\rho^{2}$ is the noise variance. Note that $\boldsymbol{K}_{\!\boldsymbol{x}_{j}\boldsymbol{x}_{j}}$ is a scalar and $\boldsymbol{K}_{\!X\boldsymbol{x}_{j}}$ is a row vector, and hence, this iterative update can be implemented efficiently.
We remark that Equations 5 and 6 are natural extensions of previous non-adaptive active learning methods, which typically maximize some notion of “distance” between points in the batch, to the “directed” setting (Ash et al., 2020; Zanette et al., 2021; Holzmüller et al., 2023; Pacchiano et al., 2024). BaCE simultaneously maximizes “distance” between points in a batch and minimizes “distance” to points in $\mathcal{A}$.
@@ G.5 Baselines
Figure 6: Comparison of “undirected” baselines for the experiment of Figure 2. In the MNIST experiment, UnSa and Undirected ITL coincide, and we therefore only plot the latter.
In the following, we briefly describe the most commonly used “undirected” decision rules.
@@
The following heuristics computed based on the softmax distribution aim to quantify the “uncertainty” about a particular input $\boldsymbol{x}$:
@@
  • MaxMargin (Scheffer et al., 2001; Settles & Craven, 2008):
    $\boldsymbol{x}_{n}=\operatorname*{arg\,min}_{\boldsymbol{x}\in\mathcal{S}}p_{1}(\boldsymbol{x};\boldsymbol{\widehat{\theta}}_{n-1})-p_{2}(\boldsymbol{x};\boldsymbol{\widehat{\theta}}_{n-1})$
    where $p_{1}$ and $p_{2}$ are the two largest class probabilities.
  • LeastConfidence (Lewis & Gale, 1994; Settles & Craven, 2008; Hendrycks & Gimpel, 2017; Tamkin et al., 2022):
    $\boldsymbol{x}_{n}=\operatorname*{arg\,min}_{\boldsymbol{x}\in\mathcal{S}}p_{1}(\boldsymbol{x};\boldsymbol{\widehat{\theta}}_{n-1})$
    where $p_{1}$ is the largest class probability.
An alternative class of decision rules aims to select diverse batches by maximizing the distances between points. Embeddings $\boldsymbol{\phi}(\boldsymbol{x})$ induce the (Euclidean) embedding distance
$d_{\boldsymbol{\phi}}(\boldsymbol{x},\boldsymbol{x^{\prime}})\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\left\|\boldsymbol{\phi}(\boldsymbol{x})-\boldsymbol{\phi}(\boldsymbol{x^{\prime}})\right\|_{2}.$
Similarly, a kernel $k$ induces the kernel distance
$d_{k}(\boldsymbol{x},\boldsymbol{x^{\prime}})\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\sqrt{k(\boldsymbol{x},\boldsymbol{x})+k(\boldsymbol{x^{\prime}},\boldsymbol{x^{\prime}})-2k(\boldsymbol{x},\boldsymbol{x^{\prime}})}.$
It is straightforward to see that if $k(\boldsymbol{x},\boldsymbol{x^{\prime}})=\boldsymbol{\phi}(\boldsymbol{x})^{\top}\boldsymbol{\phi}(\boldsymbol{x^{\prime}})$, then embedding and kernel distances coincide, i.e., $d_{\boldsymbol{\phi}}(\boldsymbol{x},\boldsymbol{x^{\prime}})=d_{k}(\boldsymbol{x},\boldsymbol{x^{\prime}})$.
  • MaxDist (Holzmüller et al., 2023; Yu & Kim, 2010; Sener & Savarese, 2017; Geifman & El-Yaniv, 2017) constructs the batch by choosing the point with the maximum distance to the nearest previously selected point:
    $\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\min_{i<n}d(\boldsymbol{x},\boldsymbol{x}_{i})$
  • Similarly, k-means++ (Holzmüller et al., 2023) selects the batch via k-means++ seeding (Arthur et al., 2007; Ostrovsky et al., 2013). That is, the first centroid $\boldsymbol{x}_{1}$ is chosen uniformly at random and the subsequent centroids are chosen with a probability proportional to the square of the distance to the nearest previously selected centroid:
    $\mathbb{P}\left(\boldsymbol{x}_{n}=\boldsymbol{x}\right)\propto\min_{i<n}d(\boldsymbol{x},\boldsymbol{x}_{i})^{2}.$
    When using the loss gradient embeddings from Equation 26, this decision rule is known as BADGE (Ash et al., 2020).
Finally, we summarize common kernel-based decision rules.
  • Undirected ITL chooses
    $\boldsymbol{x}_{n}$ $=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\mathrm{I}\!\left(\boldsymbol{f}_{\!\!\mathcal{S}};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right)$
    $=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\mathrm{I}\!\left(f_{\boldsymbol{x}};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right).$
    This can be shown to be equivalent to MaxDet (Holzmüller et al., 2023) which selects
    $\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\left|\boldsymbol{K}_{\!\boldsymbol{x}}+\sigma^{2}\boldsymbol{I}\right|$
    where $\boldsymbol{K}_{\!\boldsymbol{x}}$ denotes the kernel matrix over $\boldsymbol{x}_{1:n-1}\cup\{\boldsymbol{x}\}$, conditioned on observations $\mathcal{D}_{n-1}$.
  • Uncertainty sampling (UnSa, Lewis & Catlett, 1994) which with embeddings $\boldsymbol{\phi}_{n-1}$ after round $n-1$ selects:
@@
    With batch size $b=1$, UnSa coincides with Undirected ITL. When evaluated with gradient embeddings, this acquisition function is similar to previously used “embedding length” or “gradient length” heuristics (Settles & Craven, 2008).
We compare to the abovementioned decision rules and summarize the results in Figure 6. We observe that most “undirected” decision rules perform worse (and often significantly so) than Random. This is likely due to frequently selecting points from the support of $\mathcal{P}_{\!\mathcal{S}}$ which are not in the support of $\mathcal{P}_{\!\!\mathcal{A}}$ since the points are “adversarial examples” that the model $\boldsymbol{\widehat{\theta}}$ is not trained to perform well on. In the case of MNIST, the poor performance can also partially be attributed to the well-known “cold-start problem” (Gao et al., 2020).
In Figure 2, we also compare to the following “directed” decision rules:
@@
    $\angle_{\boldsymbol{\phi}}(\boldsymbol{x},\mathcal{A})\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\frac{1}{|\mathcal{A}|}\sum_{\boldsymbol{x^{\prime}}\in\mathcal{A}}\frac{\boldsymbol{\phi}(\boldsymbol{x})^{\top}\boldsymbol{\phi}(\boldsymbol{x^{\prime}})}{\left\|\boldsymbol{\phi}(\boldsymbol{x})\right\|_{2}\left\|\boldsymbol{\phi}(\boldsymbol{x^{\prime}})\right\|_{2}}.$
@@
    $\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\mathrm{H}[p(\boldsymbol{x};\boldsymbol{\widehat{\theta}}_{n-1})]\cdot\left(\angle_{\boldsymbol{\phi}_{n-1}}(\boldsymbol{x},\mathcal{A})\right)^{\beta}$